# Kinematic censorship and high energy particle collisions in the Schwarzschild background

### v2

@@ VIII Particle decay near singularity
IX Summary and conclusions
High energy collisions in the Schwarzschild metric are possible due to participation of a particle coming from a white hole region that are absent in the standard BSW effect. The main properties of such collisions are summarized in Table I. We indicated in scenario 1 proximity of particle 1 to the bifurcation point. However, the reservation is in order here that this is ”virtual” proximity for the trajectory which would continue without collision with particle 2. In line 2 we indicate the Schwarzschild time of collision $t_{c}$. We would like to stress that the signs $\infty$- ∞ or +$\infty$+ ∞ are pure conditional and used for the sake of brevity. They mean that in the horizon limit $r_{c}\rightarrow r_{+}$ the corresponding quantity tends to plus or minus infinity, although for any $r_{c}>r_{+}$  it is finite.
Scenario 1 Scenario 2
$\hat{t}_{c}$ finite +$\infty$+ ∞
$t_{c}$ finite +$\infty$+ ∞
$t_{0}^{(1)}$$\infty$- ∞ finite
$t_{0}^{(2)}$ finite +$\infty$+ ∞
proximity to bifurcation point particle 1 (virtually) particle 2

### v3

@@ IX Particle decay near singularity
X Summary and conclusions
High energy collisions in the Schwarzschild metric are possible due to participation of a particle coming from a white hole region that are absent in the standard BSW effect. The main properties of such collisions are summarized in Table I. We indicated in scenario 1 proximity of particle 1 to the bifurcation point. However, the reservation is in order here that this is ”virtual” proximity for the trajectory which would continue without collision with particle 2. In lines 1 and 2 we indicate the time of collision. We would like to stress that the symbols $\infty$- ∞ or +$\infty$+ ∞ are pure conditional and used for the sake of brevity. They mean that in the horizon limit $r_{c}\rightarrow r_{+}$ the corresponding quantity tends to plus or minus infinity, although for any $r_{c}>r_{+}$  it is finite. Scenario 3 combines features of those 1 and 2.
Scenario 1 Scenario 2 Scenario 3
$t_{c}$ finite +$\infty$+ ∞ +$\infty$+ ∞
$\hat{t}_{c}$ finite +$\infty$+ ∞ +$\infty$+ ∞
$t_{0}^{(1)}$$\infty$- ∞ finite $\infty$- ∞
$t_{0}^{(2)}$ finite +$\infty$+ ∞ +$\infty$+ ∞
proximity to bifurcation point particle 1 (virtually) particle 2 both